please help solve this question
The soccer coach needs new soccer balls he noticed that 6 out of the 10 soccer balls were old. he ordered new ones and has told them that soon 80% of their soccer will be new. How many new soccer balls did he order?

Answer by richwmiller(17219)

4/10 are new
Is he replacing old ones with new ones or just adding new ones?
If he replaced old with new then he bought 4 new ones and threw out 4 old old ones leaving two old ones and 8 new ones.
If he kept the old ones,
(4+x)/(10+x)=8/10
40+10x=80+8x
2x=40
x=20
So if he kept the old ones he bought 20 new ones
check
(4+20)/(10+20)=24/30=8/10
6 of the 30 are still old